# Learning hidden cascades via classification

### v2

@@ Abstract
The spreading dynamics in social networks are often studied under the assumption that individuals’ statuses, whether informed or infected, are fully observable. However, in many real-world situations, such statuses remain unobservable, which is crucial for determining an individual’s potential to further spread the infection. While this final status is hidden, intermediate indicators such as symptoms of infection are observable and provide important insights into the spread process. We propose a partial observability-aware Machine Learning framework to learn the characteristics of the spreading model. We term the method Distribution Classification, which utilizes the power of classifiers to infer the underlying transmission dynamics. We evaluate our method on two types of synthetic networks and extend the study to a real-world insider trading network. Results show that the method performs well, especially on complex networks with high cyclic connectivity, supporting its utility in analyzing real-world spreading phenomena where direct observation of individual statuses is not possible.
Keywords: Machine Learning, Hidden Cascades, Social Networks, Classification
@@ 1 Introduction
Understanding the dynamics of spreading in networks is often challenging due to the absence of a comprehensive view of the connections between individuals involved in transmission. However, in today’s increasingly digital environment, where user interactions, transactions, and communications are routinely logged and stored across platforms, reconstructing the transmission network is becoming more feasible (ZHOU201792, ). The data may be noisy or incomplete, but the availability of large-scale digital traces offers a valuable foundation for inferring transmission pathways. Learning the transmission dynamics of contagion, whether in the context of disease, insider trading, or information spread, can be achieved by the network approach. It requires that the network structure is a meaningful substrate for the underlying transmission pathways (Dutta2018, ). When this condition is satisfied, diverse spreading phenomena can in theory be analyzed within a common analytical framework, enabling consistent modeling across different domains.
Traditional studies often assume full observability of individuals’ transmission statuses, but in practice, such visibility is rarely available (ZHOU201792, ; pouget2015inferring, ; newman2023message, ; wilinski2021prediction, ). For example, during the COVID-19 pandemic, infection chains were frequently untraceable due to asymptomatic cases, misleading symptoms, and unreliable rapid tests. Similarly, in financial markets, the spread of private information through social connections is typically unobservable, making it difficult to identify who holds privileged information. In such scenarios, it remains unclear whether individuals were the carriers (either through being informed or infected) or when the transmission occurred. The absence of temporal and status information makes conventional Maximum Likelihood Estimation (MLE) methods unsuitable gomez2012inferring . While recent research has started to address partially observed data ramezani2023joint , existing methods fail to account for both hidden infection states and indirect symptom-based observations; in this work, we address this challenge, referred to as the Hidden Cascade (HC) problem.
Our method addresses the question of hidden or unreliable node status using symptom-based indirect observations in the context of cascades, and as such is a generalization of missing node status, as the nodes may exhibit false positive or false negative symptoms, not just missing observations.
@@
Figure 1 illustrates the HCs, where the source of propagation is shown as a black node. In each cascade event, links between nodes are randomly activated with a propagation probability $p$, which represents the likelihood of propagation. The value of $p$ is not directly observable and is the main focus of inference in this work. We classify transmission links between nodes into three types: (1) successful transmissions (blue), where the target node is successfully infected by the source; (2) failed transmissions (red), where the source attempts transmission but the target does not receive it; and (3) non-transmissions (black), where the source node is not a carrier. The upper semicircle of each node indicates its status: blue represents an infected (carrier) node, while white denotes an non-carrier node. In HCs, the true status of individuals, such as whether they are carriers or not, is unobservable. Instead, observations are limited to symptoms, which are driven by the symptom probability $q$, applicable across both information and infection cascades. Positive symptoms, shown as green lower semicircles, are indicators that a node might be a carrier, but this cannot be confirmed with certainty. In information cascades, symptoms reflect behaviors influenced by the possible reception of information. For example, an investor connected to company insiders who makes a profitable trade before a public announcement, may exhibit behavior consistent with prior access to private information, although this is not definitive proof. A node can also be a carrier without displaying symptoms, referred to as an asymptomatic case (for example, node 8). On the other hand, symptoms may also be false positives, where a node shows signs that appear to indicate carrier status but are unrelated (as in node 2). Anti-symptoms, or negative symptoms, represented by red lower semicircles, may also occur. These suggest that a node is unlikely to be a carrier, though again not with certainty. Examples include an agent making a loss-making trade before a public announcement, suggesting they were likely not privately informed, or a person showing test results that contradict the usual disease profile. There is also a non-symptomatic state, where the individual shows no symptoms related to being a carrier. The overall transmission in the HCs is governed by two probabilities: the propagation probability $p$ and the symptom probability $q$.
@@
The framework proposed in this paper is motivated by the principle of distribution matching, a technique that has been successfully applied across a variety of domains. Distribution matching enables synthetic data generation by preserving the statistical or receptive field properties of the original data, supporting tasks such as dataset condensation hinton2015 ; zhao2023dataset and graph condensation liu2022graph . While distribution matching has demonstrated its versatility across various applications, this research introduces a novel framework termed Distribution Classification (DC). The proposed approach infers the underlying parameters of a spreading model $\psi(\theta)$ by classifying summary statistics, which serve as a holistic representation of the distributions from which the features are drawn. Rather than directly estimating the parameters, DC employs an adversarial strategy where a classifier is trained to be maximally uncertain in distinguishing between summary statistics generated from the ground truth parameter setting and those from sampled configurations, using this induced confusion as a mechanism for parameter inference.
@@ 2 Related Work
Epidemic Spread. The study of epidemic spread has long been a prominent area of scientific inquiry and remains highly active to this day. One of the earliest and most notable contributions dates back to the 19th century, when John Snow investigated the spread of cholera in London, laying the foundation for modern epidemiology snow2023 . Since then, extensive research has been dedicated to understanding how infectious diseases propagate through populations. In recent years, the spread of COVID-19 has been modeled probabilistically to capture its transmission dynamics kucharski2020 ; bherwani2021 ; saxena2021 . These models have supported vaccine and drug development and guided public health interventions. Classical models such as the susceptible-infected (SI) and susceptible-infected-recovered (SIR) frameworks Dutta2018 have been widely used to simulate and analyze such dynamics.
@@
Information Spread. Beyond biological contagions, similar spreading processes appear in other domains, including social and financial networks. For instance, in financial settings, investors often share valuable non-public information with close contacts to enable profitable trades. Recent studies have used topological clustering and graph neural networks to identify individuals likely to receive such illicit information goel2024 ; baltakys2023predicting , confirming the presence of hidden influence paths within these networks. These parallels between epidemiological and informational spreading highlight the broader utility of contagion modeling across domains.
@@
Existing methods. A key challenge across both domains is reconstructing the underlying transmission pathways from limited observations. Maximum Likelihood Estimation (MLE) methods rely on fully observable timestamped data gomez2012inferring , being unsuitable when only the final carrier statuses are known. To address it, prior work has explored inference strategies based on Message Passing methods, particularly the dynamic message passing (DMP) framework lokhov2016reconstructing ; lokhov2017optimal ; wilinski2021prediction ; wilinski2024learning . While DMP performs well in sparse, tree-like networks, its accuracy diminishes in dense, loopy graphs, such as those found in financial networks, where short cycles violate its assumptions. Moreover, in many real-world settings, we observe only the final symptoms at the end of a time window, without access to the full temporal trajectory or the hidden infection state, defining the Hidden Cascade (HC) problem. These limitations have motivated us to develop the Hidden Cascade inference framework, which seeks to infer transmission dynamics from partial or noisy final observations. In such settings, observations may be indirect and uncertain, yet many existing approaches still assume full observability and certainty in carrier statuses. This mismatch restricts their applicability in real-world contexts, where observations are inherently partial and noisy.
@@
In developing a framework for learning spreading processes through likelihood-free inference, we were inspired by the work of gutmann2018likelihood , which introduced a likelihood-free inference framework based on classification. Methodologically, their framework differs from ours, particularly because spreading processes on graphs require the use of distribution matching. While spreading processes have previously been inferred in a likelihood-free manner, the approaches used differ from the principles we adopt in this paper. Dutta2018 employed an approximate Bayesian computation method to simultaneously learn the parameters of the spreading process and identify the initially infected node. More recently, wang2024flexible developed a Bayesian inference approach for estimating the parameters of a partially observed contagious process. Deep learning methods have also been explored; for example, murphy2021deep used deep graph neural networks to forecast the evolution of contagion dynamics.
@@ 3.2.1 Modeling Hidden Cascades with the Independent Cascade Model
To simulate hidden cascades from the underlying infection processes, we introduce a probabilistic observation model based on symptom generation. Instead of directly observing the binary infection states of nodes, we observe noisy symptoms associated with each node. Each node $v\in V$ exhibits a symptom variable $z_{v}\in\{-1,0,+1\}$, where $+1$ denotes a positive symptom, $-1$ a negative symptom, and 00 indicates the absence of symptoms. The observed symptom $z_{v}$ is generated according to a conditional distribution based on the true (latent) infection state $a_{v}\in\{0,1\}$, where $a_{v}=1$ indicates infection and $a_{v}=0$ otherwise.
@@
As a result, the simulated data comprises symptom vectors instead of infection labels, introducing noise and ambiguity akin to real-world observations. (see Appendix A for details)
@@ 3.2.2 Feature Generation from Noisy Symptom Observations
To further model uncertainty in the diffusion process, we employ a Monte Carlo-based approach. For each node, we simulate $N$ independent cascades and compute symptom-based summary statistics. This procedure is repeated $M$ times, producing $M$ feature vectors per node. These samples collectively capture the distributional structure induced by the propagation and observation models, offering a compact yet expressive encoding of the node’s diffusion behavior. Further implementation details, including pseudocode, are provided in Appendix B (Algorithm 1).
@@ 3.2.3 Model Parameter Optimization
Due to the nature of our objective function, which is evaluated through Monte Carlo simulations, an analytical gradient is not available. The function is non-differentiable, making gradient-based optimization methods unsuitable. Consequently, we employ Powell’s conjugate direction method, a derivative-free optimization algorithm that sequentially refines a set of search directions through one-dimensional minimizations powell1964efficient . Rather than seeking a strict minimum directly, this method monitors changes in both the objective function (CA) and the parameter updates, using two tolerances: function tolerance $\mathrm{ftol}$ and parameter tolerance $\mathrm{xtol}$. Convergence is achieved when the improvement in the objective function between successive iterations falls below $\mathrm{ftol}$, and the maximum change in any parameter is less than $\mathrm{xtol}$. Powell’s method iteratively updates the search directions in the parameter space until these stopping criteria are met, ensuring stable and robust convergence. The learning procedure for optimal parameters with the associated algorithms is detailed in Appendix  B (see Algorithm 2).
@@ 3.2.4 Hyperparameter Selection for Classifier
While the goal is to train the model to exhibit maximum uncertainty in classification, it is equally important to ensure optimal performance. Hyperparameter optimization is essential to achieve the best settings for each classifier, balancing model accuracy with robustness under uncertainty. This ensures that the model not only captures data variability effectively but also performs efficiently in learning the propagation probabilities. In our approach, we train a separate classifier for each entity. Since the feature vectors depend on the transmission parameters $\theta$, every time a new $\theta$ is drawn from the parameter space $\Theta$, the feature vectors change—requiring the classifier to retune its hyperparameters. However, performing hyperparameter tuning at every iteration would be computationally expensive. To address this, we adopt the following strategy: hyperparameter tuning is performed only at the iteration where Powell’s optimization converges. At this stage, for each entity, the classifier is fine-tuned using the feature vectors derived from the converged $\theta$. If the resulting classifier achieves higher classification accuracy than before, Powell’s optimization is restarted using the previously converged $\theta$ as the new initial guess. The classifier tuning algorithm is provided in Appendix  B (see Algorithm 3). The hyperparameter search space used for all selected classifiers is provided in Appendix C.
@@ 4.1 Evaluation on Synthetic Graphs
We evaluated our method on two network structures: a balanced tree graph and a Barabási–Albert graph with $k=2$ leads to a graph with a high density of connections and the presence of loops, as each new node connects to only a few existing nodes, rapidly creating densely interconnected structures. (see Figure 3(a) and 3(b) in Appendix D ). Initially, only one entity (the seed) is the infected at time $t=0$, acting as the carrier that begins transmitting to its neighboring entities over subsequent time points. We tested multiple classifiers, including Decision Tree, K-Nearest Neighbors (KNN), Logistic Regression, Naive Bayes, Random Forest, Stochastic Gradient Descent (SGD), and Support Vector Machine (SVM). In addition to the summary statistics discussed in Section 3.2.2, we extended the feature vector by including the mean, variance, entropy, and the change in the number of positive, negative, and no-symptom cases between earlier and later time intervals, capturing temporal variation in symptom occurrence. This extension aimed to capture temporal variations in symptom progression. To rigorously assess the robustness of our approach, we systematically varied the parameters across a range of configurations and evaluated the corresponding impact on classifier performance.
@@
Tree Decision Tree 0.26 ± 0.07 0.74 ± 0.08 $6.16\times 10^{-3}$ 0.49% 0.26 ± 0.03 0.71 ± 0.09 $5.24\times 10^{-3}$ 0.48%
KNN 0.28 ± 0.05 0.77 ± 0.10 $7.61\times 10^{-3}$ 0.49% 0.22 ± 0.14 0.39 ± 0.35 $1.09\times 10^{-1}$ 0.48%
Logistic Regression 0.95 ± 0.03 0.94 ± 0.11 $2.43\times 10^{-1}$ 0.43% 0.21 ± 0.09 0.68 ± 0.04 $8.37\times 10^{-3}$ 0.46%
Naive Bayes 0.30 ± 0.03 0.72 ± 0.04 $1.10\times 10^{-3}$ 0.47% 0.25 ± 0.09 0.58 ± 0.27 $4.17\times 10^{-2}$ 0.49%
Random Forest 0.24 ± 0.08 0.73 ± 0.03 $4.99\times 10^{-3}$ 0.48% 0.30 ± 0.02 0.71 ± 0.06 $1.74\times 10^{-3}$ 0.48%
SGD 0.27 ± 0.04 0.70 ± 0.04 $1.67\times 10^{-3}$ 0.47% 0.29 ± 0.18 0.64 ± 0.09 $1.79\times 10^{-2}$ 0.47%
SVM 0.31 ± 0.01 0.67 ± 0.03 $1.04\times 10^{-3}$ 0.43% 0.28 ± 0.03 0.73 ± 0.02 $1.18\times 10^{-3}$ 0.43%
Loopy Decision Tree 0.30 ± 0.02 0.73 ± 0.09 $3.86\times 10^{-3}$ 0.51% 0.26 ± 0.09 0.68 ± 0.02 $4.21\times 10^{-3}$ 0.58%
KNN 0.22 ± 0.11 0.46 ± 0.33 $8.00\times 10^{-2}$ 0.67% 0.28 ± 0.04 0.69 ± 0.06 $2.25\times 10^{-3}$ 0.48%
Logistic Regression 0.47 ± 0.15 0.38 ± 0.00 $7.46\times 10^{-2}$ 0.43% 0.28 ± 0.01 0.73 ± 0.10 $4.46\times 10^{-3}$ 0.46%
Naive Bayes 0.22 ± 0.11 0.47 ± 0.33 $8.01\times 10^{-2}$ 0.67% 0.18 ± 0.11 0.34 ± 0.33 $1.20\times 10^{-1}$ 0.78%
Random Forest 0.25 ± 0.09 0.63 ± 0.31 $4.45\times 10^{-2}$ 0.61% 0.24 ± 0.13 0.70 ± 0.02 $7.67\times 10^{-3}$ 0.57%
SGD 0.17 ± 0.10 0.52 ± 0.39 $8.93\times 10^{-2}$ 0.75% 0.27 ± 0.03 0.75 ± 0.10 $6.03\times 10^{-3}$ 0.48%
SVM 0.30 ± 0.00 0.69 ± 0.0047 $4.36\times 10^{-5}$ 0.42% 0.30 ± 0.07 0.70 ± 0.02 $1.93\times 10^{-4}$ 0.43%
@@
From Table 1, it can be confirmed that the SVM classifier, with its optimal hyperparameter setting, outperforms the other classifiers, and that the optimal feature set proposed in Section 3.2.2 more effectively captures the underlying spreading model’s parameters; hence, we fix both the classifier and the summary statistic for all subsequent experiments. In Appendix  E, results on robustness checks with different propagation and symptom probabilities for trees and loopy graphs are available (see Tables 2 and 3). To compare the distribution of actual and simulated data under the learned parameters, we split nodes by distance from the seed and by degree. Figure 2 shows that the symptom distribution in the simulated cascades closely matches that of the real data. Nodes closer to the seed and those with higher degree tend to exhibit more positive symptoms, capturing key structural effects observed in real cascades.

### v3

@@ Abstract
The spreading dynamics in social networks are often studied under the assumption that individuals’ statuses, whether informed or infected, are fully observable. However, in many real-world situations, such statuses remain unobservable, which is crucial for determining an individual’s potential to further spread the infection. While final statuses are hidden, intermediate indicators such as symptoms of infection are observable and provide useful representations of the underlying diffusion process. We propose a partial observability-aware Machine Learning framework to learn the characteristics of the spreading model. We term the method Distribution Classification, which utilizes the power of classifiers to infer the underlying transmission dynamics. Through extensive benchmarking against Approximate Bayesian Computation and GNN-based baselines, our framework consistently outperforms these state-of-the-art methods, delivering accurate parameter estimates across diverse diffusion settings while scaling efficiently to large networks. We validate the method on synthetic networks and extend the study to a real-world insider trading network, demonstrating its effectiveness in analyzing spreading phenomena where direct observation of individual statuses is not possible.
@@ 1 Introduction
Understanding the dynamics of spreading in networks is often challenging due to the absence of a comprehensive view of the connections between individuals involved in transmission. However, in today’s increasingly digital environment, where user interactions, transactions, and communications are routinely logged and stored across platforms, reconstructing the transmission network is becoming more feasible (Zhou et al., 2017). The data may be noisy or incomplete, but the availability of large-scale digital traces offers a valuable foundation for inferring transmission pathways. Learning the transmission dynamics of contagion, whether in the context of disease, insider trading, or information spread, can be achieved by the network approach. It requires that the network structure is a meaningful substrate for the underlying transmission pathways (Dutta et al., 2018). When this condition is satisfied, diverse spreading phenomena can in theory be analyzed within a common analytical framework, enabling consistent modeling across different domains.
Traditional studies often assume full observability of individuals’ transmission statuses, but in practice, such visibility is rarely available (Zhou et al., 2017; Pouget-Abadie & Horel, 2015; Newman, 2023; Wilinski & Lokhov, 2021). For example, during the COVID-19 pandemic, infection chains were frequently untraceable due to asymptomatic cases, misleading symptoms, and unreliable rapid tests. Similarly, in financial markets, the spread of private information through social connections is typically unobservable, making it difficult to identify who holds privileged information. In such scenarios, it remains unclear whether individuals were the carriers (either through being informed or infected) or when the transmission occurred. The absence of temporal and status information makes conventional Maximum Likelihood Estimation (MLE) methods unsuitable Gomez-Rodriguez et al. (2012). While recent research has started to address partially observed data Ramezani et al. (2023), existing methods fail to account for both hidden infection states and indirect symptom-based observations; in this work, we address this challenge, referred to as the Hidden Cascade (HC) problem.
Our method addresses the question of hidden or unreliable node status using symptom-based indirect observations in the context of cascades, and as such is a generalization of missing node status, as the nodes may exhibit false positive or false negative symptoms, not just missing observations. Hidden cascades are also related to cascade reconstruction from partial observations, a problem that has been widely studied. Existing approaches typically assume a one-sided observation model, where infected nodes may be partially observed but uninfected nodes are never observed as infected. In contrast, we propose a two-sided observation model that also accounts for false positives, which makes the reconstruction task more challenging and highlights the novelty of our approach.
@@
Figure 1 illustrates the HCs, where the source of propagation is shown as a black node. In each cascade event, links between nodes are randomly activated with a propagation probabilities $p_{i}$, which represents the likelihood of propagation. We first assume a uniform propagation probability $p$ across all nodes, and this assumption is subsequently relaxed in Appendix C.1. We classify transmission links between nodes into three types: (1) successful transmissions (blue), where the target node is successfully infected by the source; (2) failed transmissions (red), where the source attempts transmission but the target does not receive it; and (3) non-transmissions (black), where the source node is not a carrier. The upper semicircle of each node indicates its status: blue represents an infected (carrier) node, while white denotes an non-carrier node. In HCs, the true status of individuals, such as whether they are carriers or not, is unobservable. Instead, observations are limited to symptoms, which are driven by the symptom probability $q$, applicable across both information and infection cascades. For tractability, our framework considers a uniform $q$ across all nodes, since allowing it to vary would substantially increase the complexity of estimation without contributing to the central focus of our study Gutmann et al. (2018). Positive symptoms, shown as green lower semicircles, are indicators that a node might be a carrier, but this cannot be confirmed with certainty. In information cascades, symptoms reflect behaviors influenced by the possible reception of information. For example, an investor connected to company insiders who makes a profitable trade before a public announcement, may exhibit behavior consistent with prior access to private information, although this is not definitive proof. A node can also be a carrier without displaying symptoms, referred to as an asymptomatic case (for example, node 8). On the other hand, symptoms may also be false positives, where a node shows signs that appear to indicate carrier status but are unrelated (as in node 2). Anti-symptoms, or negative symptoms, represented by red lower semicircles, may also occur. These suggest that a node is unlikely to be a carrier, though again not with certainty. Examples include an agent making a loss-making trade before a public announcement, suggesting they were likely not privately informed, or a person showing test results that contradict the usual disease profile. There is also a non-symptomatic state, where the individual shows no symptoms related to being a carrier. The overall transmission in the HCs is governed by two probabilities: the propagation probability $p$ and the symptom probability $q$.
@@
The framework proposed in this paper is motivated by the principle of distribution matching, a technique that has been successfully applied across a variety of domains. Distribution matching enables synthetic data generation by preserving the statistical or receptive field properties of the original data, supporting tasks such as dataset condensation Hinton et al. (2015); Zhao & Bilen (2023) and graph condensation Liu et al. (2022). While distribution matching has demonstrated its versatility across various applications, this research introduces a novel framework termed Distribution Classification (DC). The proposed approach infers the underlying parameters of a spreading model $\psi(\theta)$ by classifying summary statistics, which serve as a holistic representation of the distributions from which the features are drawn. Rather than directly estimating the parameters, DC employs an adversarial strategy where a classifier is trained to be maximally uncertain in distinguishing between summary statistics generated from the ground truth parameter setting and those from sampled configurations, using this induced confusion as a mechanism for parameter inference.
@@ 2 Related Work
Epidemic Spread. The study of epidemic spread has long been a prominent area of scientific inquiry and remains highly active to this day. One of the earliest and most notable contributions dates back to the 19th century, when John Snow investigated the spread of cholera in London, laying the foundation for modern epidemiology Snow (2023). Since then, extensive research has been dedicated to understanding how infectious diseases propagate through populations. In recent years, the spread of COVID-19 has been modeled probabilistically to capture its transmission dynamics Kucharski et al. (2020); Bherwani et al. (2021); Saxena et al. (2021). These models have supported vaccine and drug development and guided public health interventions. Classical models such as the susceptible-infected (SI) and susceptible-infected-recovered (SIR) frameworks Dutta et al. (2018) have been widely used to simulate and analyze such dynamics.
@@
Information Spread. Beyond biological contagions, similar spreading processes appear in other domains, including social and financial networks. For instance, in financial settings, investors often share valuable non-public information with close contacts to enable profitable trades. Recent studies have used topological clustering and graph neural networks to identify individuals likely to receive such illicit information Goel et al. (2024); Baltakys et al. (2023), confirming the presence of hidden influence paths within these networks. These parallels between epidemiological and informational spreading highlight the broader utility of contagion modeling across domains.
@@
Existing methods. A key challenge across both domains is reconstructing the underlying transmission pathways from limited observations. Maximum Likelihood Estimation (MLE) methods rely on fully observable timestamped data Gomez-Rodriguez et al. (2012), being unsuitable when only the final carrier statuses are known. To address it, prior work has explored inference strategies based on Message Passing methods, particularly the dynamic message passing (DMP) framework Lokhov (2016); Lokhov & Saad (2017); Wilinski & Lokhov (2021; 2024). While DMP performs well in sparse, tree-like networks, its accuracy diminishes in dense, loopy graphs, such as those found in financial networks, where short cycles violate its assumptions. Moreover, in many real-world settings, we observe only the final symptoms at the end of a time window, without access to the full temporal trajectory or the hidden infection state, defining the Hidden Cascade (HC) problem. These limitations have motivated us to develop the Hidden Cascade inference framework, which seeks to infer transmission dynamics from partial or noisy final observations. In such settings, observations may be indirect and uncertain, yet many existing approaches still assume full observability and certainty in carrier statuses. This mismatch restricts their applicability in real-world contexts, where observations are inherently partial and noisy.
@@
In developing a framework for learning spreading processes through likelihood-free inference, we were inspired by the work of Gutmann et al. (2018), which introduced a likelihood-free inference framework based on classification. Methodologically, their framework differs from ours, particularly because spreading processes on graphs require the use of distribution matching. While spreading processes have previously been inferred in a likelihood-free manner, the approaches used differ from the principles we adopt in this paper. Dutta et al. (2018) employed an approximate Bayesian computation method to simultaneously learn the parameters of the spreading process and identify the initially infected node. More recently, Wang & Onnela (2024) developed a Bayesian inference approach for estimating the parameters of a partially observed contagious process. Deep learning methods have also been explored; for example, Murphy et al. (2021) used deep Graph Neural Networks (GNN) to forecast the evolution of contagion dynamics.
@@ 3.2.1 Modeling Hidden Cascades with the Independent Cascade Model
To simulate hidden cascades from the underlying infection processes, we introduce a probabilistic observation model based on symptom generation. Instead of directly observing the binary infection states of nodes, we observe noisy symptoms associated with each node. Each node $v\in V$ exhibits a symptom variable $z_{v}\in\{-1,0,+1\}$, where +1+1 denotes a positive symptom, 1-1 a negative symptom, and 0 indicates the absence of symptoms. The observed symptom $z_{v}$ is generated according to a conditional distribution based on the true (latent) infection state $a_{v}\in\{0,1\}$, where $a_{v}=1$ indicates infection and $a_{v}=0$ otherwise.
@@
As a result, the simulated data comprises symptom vectors instead of infection labels, introducing noise and ambiguity akin to real-world observations (see Appendix A for details).
@@ 3.2.2 Feature Generation from Noisy Symptom Observations
To further model uncertainty in the diffusion process, we employ a Monte Carlo-based approach. For each node, we simulate $N$ independent cascades and compute symptom-based summary statistics. This procedure is repeated $M$ times, producing $M$ feature vectors per node. These samples collectively capture the distributional structure induced by the propagation and observation models, offering a compact yet expressive encoding of the node’s diffusion behavior. Further implementation details, including pseudocode, are provided in Appendix D (Algorithm 1).
@@ 3.2.3 Model Parameter Optimization
Due to the nature of our objective function, which is evaluated through Monte Carlo simulations, an analytical gradient is not available. The function is non-differentiable, making gradient-based optimization methods unsuitable. Consequently, we employ Powell’s conjugate direction method, a derivative-free optimization algorithm that sequentially refines a set of search directions through one-dimensional minimizations Powell (1964). Rather than seeking a strict minimum directly, this method monitors changes in both the objective function (CA) and the parameter updates, using two tolerances: function tolerance $\mathrm{ftol}$ and parameter tolerance $\mathrm{xtol}$. Convergence is achieved when the improvement in the objective function between successive iterations falls below $\mathrm{ftol}$, and the maximum change in any parameter is less than $\mathrm{xtol}$. Powell’s method iteratively updates the search directions in the parameter space until these stopping criteria are met, ensuring stable and robust convergence. The learning procedure for optimal parameters with the associated algorithms is detailed in Appendix  D (see Algorithm 2).
@@ 3.2.4 Hyperparameter Selection for Classifier
While the goal is to train the model to exhibit maximum uncertainty in classification, it is equally important to ensure optimal performance. Hyperparameter optimization is essential to achieve the best settings for each classifier, balancing model accuracy with robustness under uncertainty. This ensures that the model not only captures data variability effectively but also performs efficiently in learning the propagation probabilities. In our approach, we train a separate classifier for each entity. Since the feature vectors depend on the transmission parameters $\theta$, every time a new $\theta$ is drawn from the parameter space $\Theta$, the feature vectors change—requiring the classifier to retune its hyperparameters. However, performing hyperparameter tuning at every iteration would be computationally expensive. To address this, we adopt the following strategy: hyperparameter tuning is performed only at the iteration where Powell’s optimization converges. At this stage, for each entity, the classifier is fine-tuned using the feature vectors derived from the converged $\theta$. If the resulting classifier achieves higher classification accuracy than before, Powell’s optimization is restarted using the previously converged $\theta$ as the new initial guess. The classifier tuning algorithm is provided in Appendix  D (see Algorithm 3). The hyperparameter search space used for all selected classifiers is provided in Appendix E.
@@ 4.1 Evaluation on Synthetic Graphs
We evaluated our method on two network structures: a balanced tree graph and a Barabási–Albert graph with k=2k=2444k=2k=2 leads to a graph with a high density of connections and the presence of loops, as each new node connects to only a few existing nodes, rapidly creating densely interconnected structures. (see Figure 3(a) and 3(b) in Appendix F ). Initially, only one entity (the seed) is the infected at time $t=0$, acting as the carrier that begins transmitting to its neighboring entities over subsequent time points. We tested multiple classifiers, including Decision Tree, K-Nearest Neighbors (KNN), Logistic Regression, Naive Bayes, Random Forest, Stochastic Gradient Descent (SGD), and Support Vector Machine (SVM). In addition to the summary statistics discussed in Section 3.2.2, which are considered an optimal feature set, we extend this set by including the mean, variance, entropy, and the changes in the number of positive, negative, and no-symptom cases between earlier and later time intervals, thereby capturing temporal variation in symptom occurrence. This extension aimed to capture temporal variations in symptom progression. From Table 1, it can be confirmed that the SVM classifier, with its optimal hyperparameter setting, outperforms the other classifiers, and that the optimal feature set more effectively captures the underlying spreading model’s parameters; hence, we fix both the classifier and the summary statistic for all subsequent experiments. To compare the distribution of actual and simulated data under the learned parameters, we split nodes by distance from the seed and by degree. Figure 2 shows that the symptom distribution in the simulated cascades closely matches that of the real data. Nodes closer to the seed and those with higher degree tend to exhibit more positive symptoms, capturing key structural effects observed in real cascades.
@@
Tree Decision Tree 0.26 ± 0.07 0.74 ± 0.08 $6.16\times 10^{-3}$ 0.49 0.26 ± 0.03 0.71 ± 0.09 $5.24\times 10^{-3}$ 0.48
KNN 0.28 ± 0.05 0.77 ± 0.10 $7.61\times 10^{-3}$ 0.49 0.22 ± 0.14 0.39 ± 0.35 $1.09\times 10^{-1}$ 0.48
Logistic Regression 0.95 ± 0.03 0.94 ± 0.11 $2.43\times 10^{-1}$ 0.43 0.21 ± 0.09 0.68 ± 0.04 $8.37\times 10^{-3}$ 0.46
Naive Bayes 0.30 ± 0.03 0.72 ± 0.04 $1.10\times 10^{-3}$ 0.47 0.25 ± 0.09 0.58 ± 0.27 $4.17\times 10^{-2}$ 0.49
Random Forest 0.24 ± 0.08 0.73 ± 0.03 $4.99\times 10^{-3}$ 0.48 0.30 ± 0.02 0.71 ± 0.06 $1.74\times 10^{-3}$ 0.48
SGD 0.27 ± 0.04 0.70 ± 0.04 $1.67\times 10^{-3}$ 0.47 0.29 ± 0.18 0.64 ± 0.09 $1.79\times 10^{-2}$ 0.47
SVM 0.31 ± 0.01 0.67 ± 0.03 $1.04\times 10^{-3}$ 0.43 0.28 ± 0.03 0.73 ± 0.02 $1.18\times 10^{-3}$ 0.43
Loopy Decision Tree 0.30 ± 0.02 0.73 ± 0.09 $3.86\times 10^{-3}$ 0.51 0.26 ± 0.09 0.68 ± 0.02 $4.21\times 10^{-3}$ 0.58
KNN 0.22 ± 0.11 0.46 ± 0.33 $8.00\times 10^{-2}$ 0.67 0.28 ± 0.04 0.69 ± 0.06 $2.25\times 10^{-3}$ 0.48
Logistic Regression 0.47 ± 0.15 0.38 ± 0.00 $7.46\times 10^{-2}$ 0.43 0.28 ± 0.01 0.73 ± 0.10 $4.46\times 10^{-3}$ 0.46
Naive Bayes 0.22 ± 0.11 0.47 ± 0.33 $8.01\times 10^{-2}$ 0.67 0.18 ± 0.11 0.34 ± 0.33 $1.20\times 10^{-1}$ 0.78
Random Forest 0.25 ± 0.09 0.63 ± 0.31 $4.45\times 10^{-2}$ 0.61 0.24 ± 0.13 0.70 ± 0.02 $7.67\times 10^{-3}$ 0.57
SGD 0.17 ± 0.10 0.52 ± 0.39 $8.93\times 10^{-2}$ 0.75 0.27 ± 0.03 0.75 ± 0.10 $6.03\times 10^{-3}$ 0.48
SVM 0.30 ± 0.00 0.69 ± 0.0047 $4.36\times 10^{-5}$ 0.42 0.30 ± 0.07 0.70 ± 0.02 $1.93\times 10^{-4}$ 0.43